OpenStudy (lgbasallote):

how to do trigonometric substitution? could you also demonstrate examples :DDD @apoorvk

OpenStudy (apoorvk):

how to integrate? differentiate?

OpenStudy (lgbasallote):

integrate lol

OpenStudy (apoorvk):

or how to make?

OpenStudy (lgbasallote):

i dont think they have trigonometric substitution in differential?

OpenStudy (blockcolder):

Example: \[\int\frac{\sqrt{1-x^2}}{x^5}\ \text{d}x\]

OpenStudy (lgbasallote):

yes?

OpenStudy (apoorvk):

substitutions are possible in every sphere. :p very well.. let me get you a suitable example. Hmm. Lets take \[\int\sqrt{\frac{x}{{a-x}}}.dx\] So in here, the prime slogan is "Remove the Radical Sign!" \[\text{So, I look at the function, and think, hmm, if I can get an }asin^2{\theta}\] that would make it easy, so I get some form of \[\sqrt{\frac {asin^2\theta}{a - asin^2\theta}}\] Hmm okay. But what happens to the 'dx' thingy? hmm, cool \[x = asin^2\theta\] \[\text{So, } dx = 2asin\theta.cos\theta.d\theta\]

OpenStudy (apoorvk):

Now the whole integral turns out to: \[2a\int sin\theta cos\theta\sqrt{tan^2\theta}.d \theta\] \[=2a \int sin\theta cos\theta \frac {sin\theta}{cos\theta}.d\theta\] \[=\int sin^2\theta.d\theta\] now you know, \[cos(2\theta) = 1-2sin^2\theta\] \[or, sin^2\theta = \frac{1- cos2\theta}{2}\] SO the complete integral turns out to --> \[\int \frac{1-cos2\theta}{2}.d\theta\] Which I believe is pretty easily integral using what we call 'loving' integrand formulas :P

OpenStudy (apoorvk):

Now there are basic fundas, of what I can do when I notice an integral of the following types: \[\sqrt{a^2 - x^2} ---> put\text{ }in\text{ } x = asin\theta \text{ }or\text{ }acos\theta\] \[\sqrt{a^2 + x^2} ---> put\text{ }in\text{ } x = atan\theta \text{ }or\text{ } acot\theta\] \[\sqrt{ x^2- a^2} ---> put\text{ }in\text{ } x = asec\theta \text{ }or\text{ }acosec\theta\] \[\sqrt{\frac{a-x}{a+x}} --->put \text{ }in\text{ } x= acos2\theta\] See if you can visualise why we\re doing this. Any guesses? :p

OpenStudy (apoorvk):

Get this, Mr. @lgbasallote ??

OpenStudy (lgbasallote):

frankly...no :P

OpenStudy (lgbasallote):

i mean..how do you know what to substitute @apoorvk ??

OpenStudy (lgbasallote):

i also don't know where you got \(\sqrt{\tan^2 \theta}\)

OpenStudy (lgbasallote):

i also don't know how it became \(\int \sin ^2 \theta\) and onwards i don't know a lot i know =))

OpenStudy (callisto):

\[\frac{asin^2x}{a-asin^2x} =\frac{asin^2x}{a(1-sin^2x)}=\frac{sin^2x}{(1-sin^2x)}\]\[=\frac{sin^2x}{(cos^2x)}=tan^2x\]

OpenStudy (lgbasallote):

oh..why didnt i see that o.O

OpenStudy (apoorvk):